# Evaluation of Dataframe Libraries for Data Preparation on a Single Machine

### v1

@@ Abstract
In this paper, we evaluate the most popular Python dataframe libraries in general data preparation use cases to assess how they perform on a single machine. To do so, we employ 4 real-world datasets and pipelines with distinct characteristics, covering a variety of scenarios. The insights gained with this experimentation are useful to data scientists who need to choose which of the dataframe libraries best suits their data preparation task at hand. In a nutshell, we found that: for small datasets, Pandas consistently proves to be the best choice with the richest API; when RAM is limited and there is no need to complete compatibility with Pandas API, Polars is the go-to choice thanks to its resource and query optimization; when a GPU is available, CuDF often yields the best performance, while for very large datasets that cannot fit in the GPU memory and RAM, PySpark (thanks to a multi-thread execution and a query optimizer) and Vaex (exploiting a columnar data format) are the best options.
Index Terms:
data preparation, dataframe, scalability
I Introduction
@@
Companies and organizations significantly depend on their data to drive informed decisions, such as business strategy definition, supply chain management, etc. Thus, guaranteeing high data quality is fundamental for ensuring the reliability of analysis and avoiding undesired additional costs [1]. Oftentimes data is heterogeneous, i.e., data sources adopt different formats and conventions (e.g., different encodings, different ways to represent dates or numeric values, etc.), and are affected by quality issues, such as the presence of missing or duplicate values [2]. Due to that, data scientists dedicate considerable time and resources to perform data preparation [3]. Typically, this fundamental process involves multiple operations called preparators [4], combined into a pipeline [5], which aim at exploring [6], cleaning [7], and transforming raw data into curated datasets [8, 9, 10].
@@
It is said that data scientists spend up to 80% of their time on data preparation [11]. This also depends on the lack of support in the choice among many different available libraries and in the design of the pipeline that best suits both the dataset and the task at hand. The growing need for standardization and best practices has driven an increasing number of data practitioners to embrace the dataframe as the foundational data structure for constructing pipelines and developing libraries. At a high level, the dataframe is a two-dimensional data structure that consists of rows and columns [12]. Libraries for data preparation typically offer a flexible functional interface that allows to conveniently compose pipelines.
@@
In particular, Pandas [13] is by far the most widely adopted library for manipulating dataframes, and is considered by many as the de facto standard for all preparators [12]. This library has become so popular in the data science ecosystem that even the popularity of Python has been attributed sometimes to its adoption [14]. Despite its success, Pandas notoriously presents some severe limits [15]. Firstly, it has not been designed to work with large datasets efficiently. In fact, Pandas operates in a single-threaded manner, lacks support for cluster deployment, and does not implement any memory optimization mechanism (the entire data is kept in main memory until the end of the pipeline execution). To overcome these limitations, several alternative libraries have been developed and are gaining popularity among data scientists. However, these libraries come with heterogeneous features (e.g., lazy evaluation, GPU support, etc.), making it hard for data scientists to navigate among them and choose the most suitable solution.
@@
A major reason for this confusion is the lack of extensive and rigorous studies to evaluate the performance of such libraries in the data preparation scenario. Public wisdom about dataframe libraries is mostly scattered across several not peer-reviewed sources doing mutual comparisons [16, 17, 18, 19, 20] (whose major claims are mostly confirmed by our evaluation), while related examples in scientific literature [21, 22, 23] either focus on orthogonal aspects or only cover small subsets of libraries and preparators. Similarly, relational benchmarking systems such as TPC-H [24] offer limited insights about data preparation workflows [22].
TABLE I: Features of the compared dataframe libraries.
@@
Considered version 1.5.1-2.0.0 3.4.1 0.16.2 0.15.1 22.12.0 4.16.0 1.0.0
Our contribution
It is common for data scientists to perform countless iterations on samples of the datasets on their laptop/PC to define a proper data preparation pipeline for the task at hand. Hence, the faster the library employed for the pipeline execution, the less idle time they have. To support data scientists in the choice of the dataframe library that can speedup their workflow, we present the first extensive experimental comparison of the most popular alternative libraries to Pandas, namely PySpark, Modin (with Dask and Ray engines), Polars, CuDF, Vaex, and DataTable, including in the evaluation also Pandas itself and its evolution (Pandas2). For the sake of reproducibility we developed Bento, a framework that provides a general interface to design and deploy the data preparation pipeline with any library. Bento  [^1] https://github.com/dbmodena/bento allows the definition of a common Pandas-like API for all libraries, and enables the execution in a containerized environment using Docker. This process is streamlined with the help of a convenient configuration file.
To cover a variety of scenarios and to provide results that are representative of real-world use cases, we consider four datasets with heterogeneous size and features, collected from Kaggle [25] (the most popular platforms for data science challenges) and previously adopted in related work about data preparation [4] and dataframe libraries [21]. The choice of these datasets from Kaggle allows us to exploit published data preparation pipelines that have proven to work well for downstream machine learning tasks and that have been validated by the community.
We identify four essential data preparation stages in the pipelines: input/output (I/O), exploratory data analysis (EDA), data transformation (DT), and data cleaning (DC). Thus, we evaluate the performance of the different libraries both on a single preparator and on sequences of preparators (i.e., the entire pipeline or its subsets corresponding to the four different stages), highlighting the impact of the query optimization techniques supported by some of them.
For whom is this paper relevant
We focus only on the single-machine scenario. Our findings are useful to those data scientists that: (i) are employing Pandas in their workflow when devising data preparation pipelines on their laptop/pc; (ii) want to replace Pandas with a more efficient library that has the same interface. Thus, we evaluated scalability by measuring the performance of the dataframe libraries on three distinct machine configurations (laptop, workstation, and server), while increasing the size of the dataset. Hence, we provide the readers with several insights about their performance based on multiple factors, such as the size and the features of the dataset, the configuration of the underlying machine, and the type of applied preparators.
Outline
We describe in detail the dataframe data structure and the libraries to be compared in Section II, then we present our evaluation setup and the results in Section III and IV respectvely. We review the related literature in Section V, then we report the key takeaways in Section VI. In Section VII we draw the conclusions.
II Dataframes
A dataframe [12] is a data structure that organizes the data into a two-dimensional table-like structure, where columns represent the schema and rows represent the content. Each column has a specific data type, and the schema does not need to be declared in advance. A dataframe provides a wide set of operators acting on all data dimensions, enabling and simplifying various data preparation tasks such as data analysis, transformation, and cleaning. Beyond this variety of features, dataframes became popular for data analysis and manipulation thanks to their simple and flexible structure, which allows them to effectively handle different types of data. In fact, compared to using specific data preparation tools [4], dataframes offer greater simplicity and customization, enabling users to manipulate the data by writing Python code without having the effort of learning new tools.
Pandas [26] has been the first widely-adopted Python library implementing dataframes (inspired by the corresponding data structure from the R language [13]) and provides efficient and expressive data structures optimized to work with structured datasets. Pandas has established itself over time as the standard library for data manipulation and analysis. The main reasons behind its widespread popularity reside in its robust functionality and user-friendly design, making it the go-to choice for handling datasets in Python. Given its enormous diffusion, Pandas is a very mature library and the existence of a huge amount of documentation and courses make its learning by data scientists and practitioners extremely straightforward.
Nevertheless, Pandas notoriously presents several limitations [15], which become especially evident when dealing with large-scale datasets, as it lacks the optimizations required to process large amounts of data in an efficient way. For instance, Pandas does not support multiprocessing and parallel computing, and it is designed to work with in-memory datasets, lacking support for datasets that exceed memory limits. Further, it needs to materialize the intermediate result of every operation (this execution strategy is known as eager evaluation), exposing it to out-of-memory risks and preventing it from applying query optimization techniques.
While Pandas initially relied on NumPy [27] data structures for memory management, it moved to Apache Arrow [28] starting from Pandas2 [29]. Arrow is a language-independent columnar memory format for flat and hierarchical data, organized to optimize analytic operations on modern hardware, allowing Pandas2 to handle large datasets more efficiently.
To address the aforementioned limitations, many solutions have emerged to seamlessly replace Pandas in Python workflows, aiming for more efficient data processing. In the following, we describe several of these popular Python libraries that can serve as replacements for Pandas. The selection of the libraries to include in our evaluation was based on three main characteristics they must have: (i) based on the dataframe data structure; (ii) compatible with the Pandas API, to enable a fair and straightforward evaluation of each operation (note that many of these libraries, recognizing Pandas as the current reference standard, already aim at Pandas-compatibility for their API [30] or describe the parallelism between their API and the Pandas API [31]); (iii) have gained more than 1k stars on GitHub, denoting the existence of a significant user base and community support. These libraries are: Pandas [26], PySpark [32], Modin [33], Polars [34], CuDF [35], Vaex [36], and DataTable [37]. We compare the selected libraries highlighting their strengths and weaknesses, to facilitate efficient data processing and provide support to data scientists and practitioners for efficiently handling their data preparation workflows.
@@ II Dataframes
TABLE II: Compatibility of dataframe libraries with Pandas API. (✓✓) fully matches Pandas interface; (✓) different interface; ($\circ$) missing from the API, but implemented by us to the best of our efforts.
@@
Dataframe Libraries
  • Resource optimization: implementing strategies to efficiently manage memory, reducing the impact on the available resources, thus improving the overall performance.
  • Lazy evaluation: enhancing the overall efficiency by maintaining a logical plan of the operations, triggering their execution only when a specific output operation is invoked, to allow applying query optimization techniques.
  • Deploy on cluster: enabling the distribution of dataframe processing across a cluster of machines, leveraging parallel computing to enhance the overall performance (as stated in Section I, we aim to include distributed computing in our evaluation in future work).
Table I lists for each dataframe library which of these features it implements, along with other characteristics about its implementation and the version considered here.
PySpark [32] is the Python API for Apache Spark, which is primarily implemented in Scala. Spark provides the capability to work with distributed collections of Java or Python objects partitioned across a cluster, known as Relational Distributed Datasets (RDDs) [38]. PySpark dataframes are RDDs consisting of Row objects: they can be seen as relational tables, on which relational operators can be applied. Spark has been popularized by its capability to process large-scale datasets on a cluster of machines (represented within Spark as a SparkContext object). Yet, standalone mode (i.e., on a single machine) is not only supported, but is also surprisingly fast in multiple scenarios, as shown in the experimental evaluation of Section IV. PySpark dataframes support lazy evaluation, relying on a query optimizer called Catalyst [38]. Basically, Catalyst translates the pipeline of operators applied to a dataframe into a query tree of logical operators, which is then optimized to generate the final execution plan.
PySpark supports two different APIs: (i) Spark SQL, which allows combining SQL queries with Spark programs for working with structured data, to efficiently read, write, transform, and analyze data using Python and SQL; (ii) Pandas on Spark (denoted as SparkPD and known as Koalas [39] before its inclusion in PySpark starting from Spark 3.2), which enables (by adding an index to the conventional Spark dataframe) the scaling of Pandas workloads by distributing them across multiple nodes without requiring modifications to the original Pandas code for most Pandas API functions ($\sim$80%).
Modin [33] is a Python library that provides a parallel alternative to Pandas [12]. Modin adopts the Pandas data format as the default storage layer, ensuring compatibility and usability, and parallelizes the execution of the operators across multiple cores [21]. Modin is designed to dynamically switch between different partition schemes: row-based, column-based, or block-based, depending on the operation. Each partition is then processed independently by the execution engine: Dask [40] or Ray [41].
Modin employs a set of 15 main operators, called core operators, to simplify Pandas functions and build its own Pandas-like API. When core operators cannot handle a function, Modin switches to the default to Pandas mode (incurring performance costs due to communication and Pandas single-threaded nature), then it reverts to a partitioned Modin dataframe after completion. Modin addresses flexible dataframe schemas by revising schema induction rules, allowing schema omission when not needed, optimizing queries, and minimizing overhead. The system implements opportunistic evaluation [42], enabling execution based on interactions. Incremental query construction is facilitated through intermediate dataframe results, prioritizing materialization of small results and employing predictive mechanisms. Modin captures and stores ordering information separately for on-demand ordering, and it explores key-value pairs for the efficient storage of sparse dataframes.
@@
Since Modin can run using two different engines, Dask and Ray, we consider both in the evaluation (denoted as ModinD and ModinR, respectively). Dask [43] is a Python library for distributed computing, which allows working with large distributed Pandas dataframes. Dask uses dictionaries, tuples, and callables to represent task schedules. The scheduler can operate in different modes, providing flexibility to the execution. It is designed to efficiently extend memory capacity using disk space, making it suitable for systems with limited memory. We considered the inclusion of Dask in our evaluation also as an independent library, but we found it to perform better in combination with Modin (Dask is not well-suited for a single machine), hence we do not report its results here. Furthermore, Dask covers approximately 55% of the Pandas API, while Modin around 90%.
@@
Ray [44] is a general-purpose framework for parallelizing Python code, using an in-memory distributed storage system and Arrow as data format. Unlike Dask, Ray does not have built-in primitives for partitioned data. While the two engines serve different use cases, their primary goal is similar: to optimize resource usage by changing how data is stored and Python code is executed.
@@
Polars [34] is a Python library written in Rust and built on top of arrow2 (the Rust implementation of Arrow format). The adoption of Arrow as underlying data structure confers efficient data processing capabilities to Polars, optimized through parallel execution, cache-efficient algorithms, and efficient usage of resources.
@@
Polars does not use an index, but each row is indexed by its integer position in the table. Polars has developed its own Domain Specific Language (DSL) for transforming data, which aims to be very usable and ensure human readability even for complex queries. Its core components are Expressions and Contexts. Expressions facilitate concise and efficient data transformations, while Contexts categorize evaluations into three main types: (i) filtering, (ii) grouping/aggregation, (iii) selection. Polars optimizes queries through early filters and projection pushdown. Moreover, it supports both eager and lazy execution. The lazy strategy allows to run queries in a streaming manner: instead of processing the entire data at once, they can be executed in batches, lightening the load on memory and CPU, hence allowing processing bigger (and even larger-than-memory) datasets in a faster manner.
@@
CuDF [35] is a component of the RAPIDS framework [45], developed by NVIDIA, which offers a Pandas-like API and enables to run general-purpose data science pipelines on GPUs. Written in C/C++ and CUDA, it harnesses the computational power of GPUs for accelerating data processing.
@@
CuDF is built on top of Arrow and facilitates the interchange of tabular data across GPU processes. CuDF leverages parallelization to execute operations on different parts of columns simultaneously across GPU cores. This approach takes advantage of all available GPU cores, resulting in efficient and high-performance computations. Note that CuDF executes all operations on the GPU, but it does not provide any optimization strategy for the execution of the pipelines.
@@
Vaex [36] is a Python library designed to handle extremely large tabular datasets (in particular astronomical catalogs), performing fast statistical analysis and visualization on this data. The key feature of Vaex is memory optimization, achieved through streaming algorithms, memory-mapped files, and a zero-copy policy, allowing to explore datasets exceeding the available memory, supported by its vaex-core extension module, written in C.
@@
Vaex enables efficient column-wise operations by wrapping a series of NumPy arrays as columns. Only a small subset of NumPy functions benefits of lazy evaluation (e.g., exp, cos, arcsin), storing their results as computation instructions and computing them only when needed. This is achieved through virtual columns, expressions that can be added to datasets without incurring additional memory usage. Further, Vaex also supports multithreading for faster computation on multi-core CPUs. By default, Vaex avoids unnecessary data copying and loading into memory. It can process random data subsets and export data in random order for efficient resource utilization.
@@
DataTable [37] is a Python package that relies on the Frame object, similar to Pandas dataframes or SQL tables (i.e., data arranged in a two-dimensional array with rows and columns). While it shares similarities with Pandas, DataTable emphasizes speed and support for big data.
@@
DataTable, which uses a native-C implementation for all data types, is developed to support column-oriented data storage, which optimizes data access and processing. DataTable enables memory-mapping of data on disk, allowing seamless processing of out-of-memory datasets, hence it is well-suited for large-scale data analysis tasks. DataTable incorporates multi-threaded data processing, leveraging all available cores for time-consuming operations, and offers efficient algorithms for sorting, grouping, and joining operations. To minimize unnecessary data copying, it adopts the copy-on-write technique for shared data, reducing memory overhead and improving overall performance.
TABLE III: Features of the selected datasets.

### v2

# and Reggio Emilia, Italy

@@ Abstract.
In this paper, we evaluate the most popular Python dataframe libraries in general data preparation use cases to assess how they perform on a single machine. To do so, we employ 4 real-world datasets and pipelines with heterogeneous features, covering a variety of scenarios. The insights gained with this experimentation are useful to data scientists who need to choose which of the dataframe libraries best suits their data preparation task at hand. In a nutshell, we found that: for small datasets, Pandas consistently proves to be the best choice with the richest API; when data fits in RAM and there is no need for complete compatibility with Pandas API, Polars is the go-to choice thanks to its in-memory execution and query optimizations; when a GPU is available, CuDF often yields the best performance, while for very large datasets that cannot fit in the GPU memory and RAM, PySpark (thanks to a multithread execution and a query optimizer) proves to be the best options.
copyright: acmlicensedjournalyear: 2024doi: XXXXXXX.XXXXXXXconference: the 28th International Conference on Extending Database Technology; March 25–28, 2025; Barcelona, Spainisbn: 978-1-4503-XXXX-X/18/06
1. Introduction
@@
Companies and organizations significantly depend on their data to drive informed decisions, such as business strategy definition, supply chain management, etc. Thus, guaranteeing high data quality is fundamental to ensure the reliability of analysis and avoid undesired additional costs (Haug et al., 2011). Oftentimes data is heterogeneous, i.e., data sources adopt different formats and conventions (e.g., different encodings, different ways to represent dates or numeric values, etc.), and are affected by quality issues, such as missing or duplicate values (Fan, 2015). Due to that, data scientists dedicate considerable time and resources to perform data preparation (Fernandes et al., 2023). Typically, this fundamental process involves multiple operations called preparators (Hameed and Naumann, 2020), combined into a pipeline (Furche et al., 2016), aimed at exploring (Lissandrini et al., 2018), cleaning (Ilyas and Chu, 2019), and transforming raw data into curated datasets (Terrizzano et al., 2015; Rezig et al., 2019; Geerts et al., 2020).
@@
It is said that data scientists spend up to 80% of their time on data preparation (Hellerstein et al., 2018). This also depends on the lack of support in the choice among many different available libraries and in the design of the pipeline that best suits both the dataset and the task at hand. The growing need for standardization and best practices has driven an increasing number of data practitioners to embrace the dataframe as the foundational data structure for constructing pipelines and developing libraries. At a high level, the dataframe is a two-dimensional data structure that consists of rows and columns (Petersohn et al., 2020). Libraries for data preparation typically offer a flexible functional interface that allows to conveniently compose pipelines.
@@
In particular, Pandas (McKinney, 2010) is by far the most widely adopted library for manipulating dataframes, and is considered by many as the de facto standard for all preparators (Petersohn et al., 2020; Jindal et al., 2021; Psallidas et al., 2022; Lee et al., [n. d.]). This library has become so popular in the data science ecosystem that even the popularity of Python itself has been sometimes attributed to its wide adoption (Claburn, 2017). Despite its success, Pandas notoriously presents some severe limits (Emani et al., 2024; Kläbe and Hagedorn, 2021). Firstly, it has not been designed to work with large datasets efficiently. In fact, Pandas operates in a single-threaded manner, lacks support for cluster deployment, and does not implement any memory optimization mechanism (the entire data is kept in main memory until the end of the pipeline execution). To overcome these limitations, several alternative libraries have been developed and are gaining popularity among data scientists. However, these libraries come with heterogeneous features (e.g., lazy evaluation, GPU support, etc.), making it hard for data scientists to navigate among them and choose the most suitable solution.
@@
A major reason for this confusion is the lack of extensive and rigorous studies to evaluate the performance of such libraries in the data preparation scenario. Public wisdom about dataframe libraries is mostly scattered across several not peer-reviewed sources doing mutual comparisons (Schmitt, 2020; Alexander, 2023; Pinner, 2023) (whose major claims are mostly confirmed by our evaluation), while related examples in scientific literature (Petersohn et al., 2021; Rehman and Elmore, 2022; Shanbhag and Chimalakonda, 2023) either focus on orthogonal aspects or only cover small subsets of libraries and preparators.
Our contribution
It is common for data scientists to perform countless iterations on samples of the datasets on their laptop or PC to define a proper data preparation pipeline for the task at hand. Hence, the faster the employed library is, the less idle time they have. To support data scientists in the choice of the dataframe library that can speedup their workflow, we present the first extensive experimental comparison of Pandas and its most popular dataframe-based alternatives, namely PySpark, Modin (with both Dask and Ray engine), Polars, CuDF, Vaex, and DataTable. For the sake of reproducibility we developed Bento, a framework that provides a general interface to design and deploy the data preparation pipeline with any library. Bento  [^1] https://github.com/dbmodena/bento allows the definition of a common Pandas-like API for all libraries and enables the execution in a containerized environment using Docker. This process is streamlined with the help of a convenient configuration file.
To cover a variety of scenarios and to provide results that are representative of real-world use cases, we consider four datasets with heterogeneous size and features, previously adopted in related literature on data preparation (Hameed and Naumann, 2020) and dataframe libraries (Petersohn et al., 2021). All these datasets are collected from Kaggle [^2] https://www.kaggle.com, the most popular platform for data science challenges. This allows us to exploit published data preparation pipelines that have proven to work well for downstream machine learning tasks and have been validated by the world’s largest data science community.
We identify four essential data preparation stages in the pipelines: input/output (I/O), exploratory data analysis (EDA), data transformation (DT), and data cleaning (DC). Thus, we evaluate the performance of the libraries both on a single preparator and on sequences of preparators (i.e., the entire pipeline or its subsets corresponding to the four different stages), highlighting the impact of the query optimization techniques supported by some of them.
For whom is this paper relevant
Since it represents by far the most common situation for data scientists (Emani et al., 2024), we only focus on the single-machine scenario, aiming to explore the distributed scenario in a future extension [^3] Note that we decided to include libraries conceived for distributed execution, such as PySpark or Modin, since through multithreading and optimized execution plans they can improve the performance of dataframe operations even on a single machine.. Our findings are useful to those data scientists that: (i) are employing Pandas in their workflow when devising data preparation pipelines on their laptop or PC; (ii) want to replace Pandas with a more efficient library that has the same interface. Thus, we evaluate scalability by measuring the performance of the dataframe libraries on three distinct machine configurations (laptop, workstation, and server), while increasing the size of the dataset. Hence, we provide the readers with several insights about their performance based on multiple factors, such as the size and the features of the dataset, the configuration of the underlying machine, and the type of preparators applied.
Outline
We describe in detail the dataframe data structure and the libraries to be compared in Section 2, then we present our evaluation setup and the results in Section 3 and 4, respectively. After reviewing the related literature in Section 5, we report the key takeaways in Section 6 and draw the conclusions in Section 7.
2. Dataframes
A dataframe (Petersohn et al., 2020) is a data structure that organizes the data into a two-dimensional table-like format, where columns represent the schema and rows represent the content. Each column has a specific data type, and the schema does not need to be declared in advance. A dataframe provides a wide set of operators acting on all data dimensions, enabling and simplifying various data preparation tasks such as data analysis, transformation, and cleaning. Beyond this variety of features, dataframes became popular thanks to their simple and flexible structure, which allows them to effectively handle different types of data. In fact, compared to specific data preparation tools (Hameed and Naumann, 2020), dataframes offer greater simplicity and customization, enabling users to manipulate the data by writing Python code without the effort of learning to use a new tool.
Pandas  [^4] https://github.com/pandas-dev/pandas has been the first widely-adopted Python library implementing dataframes (inspired by the corresponding data structure from the R language (McKinney, 2010)) and provides efficient and expressive data structures optimized to work with structured datasets. Pandas has established itself over time as the standard library for data manipulation and analysis. The main reasons behind its widespread popularity reside in its robust functionality and user-friendly design, making it the go-to choice for handling datasets in Python. Given its enormous diffusion, Pandas is a very mature library and the existence of a huge amount of documentation and courses makes its learning extremely straightforward.
Nevertheless, Pandas also presents several notorious limitations (McKinney, 2017), which become especially evident when dealing with large-scale datasets, as it lacks the optimizations required to process large amounts of data in an efficient way. For instance, Pandas does not support multiprocessing and parallel computing, and it is designed to work with in-memory datasets, lacking support for datasets that exceed memory limits. Further, it needs to materialize the intermediate result of every operation (this execution strategy is known as eager evaluation), exposing it to out-of-memory risks and preventing it from applying query optimization techniques.
To address such limitations, many solutions have emerged to seamlessly replace Pandas in Python workflows, aiming for more efficient data processing. The selection of the Python libraries to include in our evaluation was based on three main criteria, requiring them to: (i) be based on the dataframe data structure (excluding therefore SQL-based solutions such as DuckDB (Raasveldt and Mühleisen, 2019)); (ii) be compatible with the Pandas API, to enable a fair and straightforward evaluation of each operation (note that many of these libraries, recognizing Pandas as the de facto standard, already aim at Pandas-compatibility for their API [^5] https://modin.readthedocs.io/en/latest or describe the parallelism between their API and the Pandas API [^6] https://datatable.readthedocs.io/en/latest/manual/comparison_with_pandas.html); (iii) have gained more than 1k stars on GitHub, denoting the existence of a significant user base and community support. The selected libraries are: Pandas, PySpark, Modin, Polars, CuDF, Vaex, and DataTable.
Table 1. Features of the compared dataframe libraries.
@@ 2. Dataframes
Considered version 2.2.1 3.5.1 0.29.0 0.20.23 24.04.01 4.17.0 1.1.0
Dataframe Libraries
  • Resource optimization: implementing strategies to efficiently manage memory, reducing the impact on the available resources, hence improving the overall performance.
  • Lazy evaluation: maintaining a logical plan of the operations, triggering their execution only when a specific output operation is invoked, to apply query optimization techniques.
  • Deploy on cluster: enabling the distribution of dataframe processing across a cluster of machines, leveraging parallel computing (as stated in Section 1, we aim to consider this dimension in future work).
For each dataframe library, Table 1 lists which of these features it implements, along with information about its implementation and the version considered here.
PySpark [^7] https://github.com/apache/spark is the Python API for Apache Spark, which is primarily implemented in Scala. Spark allows to work with distributed collections of Java or Python objects partitioned across a cluster, known as Relational Distributed Datasets (RDDs) (Armbrust et al., 2015). PySpark dataframes are RDDs consisting of Row objects: they can be seen as relational tables, on which relational operators can be applied. Further, they support lazy evaluation, relying on the Catalyst optimizer (Armbrust et al., 2015). While Spark owes its popularity to the capability of processing large-scale datasets on a cluster of machines, standalone (i.e., single-machine) mode is not only supported, but also surprisingly fast in multiple scenarios, as shown in Section 4.
PySpark supports two different APIs: (i) Spark SQL, which allows combining SQL queries with Spark programs for working with structured data; (ii) Pandas on Spark (denoted as SparkPD and previously known as Koalas [^8] https://github.com/databricks/koalas), which enables (by adding an index to the conventional Spark dataframe) to distribute Pandas workloads across multiple nodes without requiring modifications to the original Pandas code for most API functions ($\sim$80%).
@@
Table 2. Compatibility of dataframe libraries with Pandas API. (✓✓) fully matches Pandas interface; (✓) different interface; ($\circ$) missing from the API, but implemented by us to the best of our efforts.
@@
Modin [^9] https://github.com/modin-project/modin is a Python library that provides a parallel alternative to Pandas (Petersohn et al., 2020, 2021). Modin adopts the Pandas data format as the default storage layer and employs a set of 15 core operators to simplify Pandas functions and build its own Pandas-like API. When core operators cannot handle a function, it switches to the default to Pandas mode (affecting performance due to communication costs and Pandas single-threaded nature), reverting to a partitioned Modin dataframe after completion. It also implements opportunistic evaluation (Xin et al., 2021), enabling execution based on interactions, and facilitates incremental query construction through intermediate dataframe results. Modin is designed to dynamically switch between different partition schemes (row-based, column-based, or block-based) depending on the operation. Each partition is then processed independently by the execution engine: Dask or Ray (we consider both solutions, denoted as ModinD and ModinR, respectively).
@@
Dask  [^10] https://github.com/dask/dask is a Python library for distributed computing, which allows working with large distributed Pandas dataframes (Rocklin, 2015). It is designed to efficiently extend memory capacity using disk space, making it suitable for systems with limited memory, and its scheduler provides flexibility to the execution. While we also considered the inclusion of Dask as an independent library, we do not report its results since we found it to perform better in combination with Modin (as Dask is not well suitable for a single machine) and it only covers $\sim$55of the Pandas API (while Modin $\sim$90%).
@@
Ray  [^11] https://github.com/ray-project/ray is a general-purpose framework for parallelizing Python code, using an in-memory distributed storage system and Apache Arrow [^12] https://github.com/apache/arrow (a language-independent columnar memory format, recently adopted even by Pandas  [^13] https://pandas.pydata.org/docs/dev/whatsnew/v2.0.0.html) as data format (Moritz et al., 2018). Unlike Dask, Ray does not have built-in primitives for partitioned data. While the two engines serve different use cases, their primary goal is similar: optimizing resource usage by changing how data is stored and Python code is executed.
@@
Polars [^14] https://github.com/pola-rs/polars is a Python library written in Rust and built on top of arrow2 (the Rust implementation of Arrow format). The adoption of Arrow as underlying data structure confers efficient data processing capabilities to Polars, optimized through parallel execution, cache-efficient algorithms, and efficient usage of resources.
@@
Polars does not use an index, but each row is indexed by its integer position in the table. Further, it has developed its own Domain Specific Language (DSL) for transforming data, whose core components are Expressions and Contexts. Expressions facilitate concise and efficient data transformations, while Contexts categorize evaluations into three main types: filtering, grouping/aggregation, and selection. Polars optimizes queries through early filters and projection pushdown. Moreover, it supports both eager and lazy execution. The lazy strategy allows to run queries in a streaming manner: instead of processing the entire data at once, they can run in batches, lightening the load on memory and CPU, hence allowing to process bigger datasets (even larger than memory) in a faster manner.
@@
CuDF [^15] https://github.com/rapidsai/cudf is a component of the NVIDIA RAPIDS [^16] https://rapids.ai framework. Written in C/C++ and CUDA (hence only compatible with NVIDIA GPUs), it offers a Pandas-like API to run general-purpose data science pipelines on GPUs, leveraging their computational power for accelerating data processing.
@@
CuDF is built on top of Arrow and leverages parallelization to execute operations on different parts of columns simultaneously across all available GPU cores. Note that CuDF uses a single GPU, while Dask-CuDF can be used for multi-GPU parallel computing. Thus, it can perform efficient and high-performance computations, although it does not provide any optimization strategy for the execution of the pipelines.
@@
Vaex [^17] https://github.com/vaexio/vaex is a Python library [^18] Since the last commit on GitHub is dated about 10 months ago, the project might not be actively maintained at the moment. designed to handle extremely large tabular datasets (such as astronomical catalogs), producing fast statistical analysis and visualization. Supported by its vaex-core extension, written in C, it achieves memory optimization through streaming algorithms, memory-mapped files, and a zero-copy policy, allowing to explore datasets exceeding the available memory.
@@
Vaex enables efficient column-wise operations by wrapping a series of NumPy arrays as columns. Only a small subset of NumPy functions benefits of lazy evaluation, storing their results as computation instructions and computing them only when needed. This is achieved through virtual columns, expressions that can be added to datasets without incurring additional memory usage. Vaex can process random data subsets and export data in random order for efficient resource utilization. Further, it also supports multithreading for faster computation on multi-core CPUs.
@@
DataTable [^19] https://github.com/h2oai/datatable is a Python package that relies on the Frame object, similar to Pandas dataframes or SQL tables (i.e., data arranged in a two-dimensional array with rows and columns).
@@
DataTable, which uses a native-C implementation for all data types, is developed to support column-oriented data storage, optimizing data access and processing. DataTable enables memory-mapping of data on disk, allowing the seamless processing of out-of-memory datasets. DataTable incorporates multithreaded data processing, leveraging all available cores for time-consuming operations, and offers efficient algorithms for sorting, grouping, and joining operations. To minimize unnecessary data copying, it adopts the copy-on-write technique for shared data, reducing memory overhead and improving the overall performance.
3. Bento
In this section, we provide all details about the configuration of Bento, our framework for evaluating different dataframe libraries on data preparation pipelines.
Datasets
For a comprehensive evaluation of the dataframe libraries on general, real-world data preparation use cases, we selected four datasets covering a variety of sizes, complexities, and features: (i) Athlete  [^20] https://www.kaggle.com/datasets/heesoo37/120-years-of-olympic-history-athletes-and-results, collecting 0.2M records about the results achieved by the athletes through 120 years of Olympics; (ii) Loan  [^21] https://www.kaggle.com/datasets/wordsforthewise/lending-club, containing 2M rows about loan applicants and their financial profiles from the LendingClub company; (iii) Patrol  [^22] https://www.kaggle.com/datasets/faressayah/stanford-open-policing-project, composed of 27M records about 11 years of traffic stops by California law enforcement agencies (Pierson et al., 2020); (iv) Taxi  [^23] https://www.kaggle.com/competitions/nyc-taxi-trip-duration, containing 77M records about taxi trips in New York City in 2015. A summary of their main features is displayed in Table 3.
Collected from Kaggle, all of these datasets had previously been adopted in related work about data preparation and dataframe libraries. The three largest datasets were indeed used to evaluate the performance of Modin (Petersohn et al., 2021), due to the variety of their features. Instead, Athlete was exploited to analyze how different preparators are supported by commercial data preparation tools (Hameed and Naumann, 2020).
Table 3. Features of the selected datasets.